# Ideal transition systems

### v2

@@ 3.2 Examples
In these examples, $A$ will be some affine variety of matrices over $\mathbb{K}$. Given $w\in A$, we let
@@
Our first example, while particularly simple, is helpful to illustrate the general setup. Suppose $A$ is the variety of skew-symmetric $3\times 3$ matrices
@@
$J_{k}=\operatorname{init}(I_{k})\quad\text{and}\quad\mathcal{T}=\{(I_{k},J_{k}):k\in[7]\}.$
We now investigate whether $\mathcal{T}$ is a transition system. One can informally express
$X_{1}$ =$=\left\{\begin{bmatrix}0&0\\ 0&0\end{bmatrix}\right\},\quad X_{2}=\left\{\begin{bmatrix}0&0\\ 0&\ast\end{bmatrix}\right\},\quad X_{3}=\left\{\begin{bmatrix}0&0\\ \ast&\ast\end{bmatrix}\right\},\quad X_{4}=\left\{\begin{bmatrix}0&\ast\\ 0&\ast\end{bmatrix}\right\},$
$X_{5}$ =$=\left\{\begin{bmatrix}0&\ast\\ \ast&\ast\end{bmatrix}\right\},\quad X_{6}=\left\{\begin{bmatrix}\ast&\ast\\ \ast&\ast\end{bmatrix}\right\},\quad X_{7}=\left\{\det=0\right\},$
so one has
@@
We only need to verify conditions (T3) and (T4) in Definition 3.1 when $(I,J)=(I_{7},J_{7})$. First suppose $J_{7}=\langle x_{12}x_{21}\rangle$. Then $I_{7}+\langle x_{11}\rangle=J_{7}+\langle x_{11}\rangle=\langle x_{11},x_{12}x_{21}\rangle=I_{3}\cap I_{4}=J_{3}\cap J_{4},$ so $f=x_{11}$ and $\Phi=\{(I_{3},J_{3}),(I_{4},J_{4})\}$ satisfy condition (T3) in Definition 3.1. For these choices, condition (T4) holds vacuously, and we conclude that $\mathcal{T}$ is a transition system.
@@
The set $\{(I_{k},J_{k}):k\in[5]\}$ fails to be a transition system, though only by a small margin.
We only need to verify conditions (T3) and (T4) in Definition 3.1 when $(I,J)=(I_{1},J_{1})$ since $I_{k}=J_{k}$ when $k>1$. The problem is that $J_{1}+\langle\operatorname{init}(f)\rangle$ is never equal to $\bigcap_{k\in S}J_{k}$ for any choice of set $\varnothing\subsetneq S\subseteq[5]$ when $f\in\mathbb{K}[A]$ is non-constant with $f\notin I_{1}$. To get around this, define
@@ 3.3 Orbit closures
We are most interested in transitions systems $\mathcal{T}$ made up of pairs $(I,J)$ where $I$ ranges over the orbit closures of certain maximal ideals. This section explains the general properties of the relevant orbit closure operation. It is convenient to present these results in an entirely self-contained way, but generalizations of this material in terms of equivariant sheaves may be well-known; see, e.g., [5, 30].
Continue to let $A$ be an affine variety over an algebraically closed field $\mathbb{K}$. Suppose $G$ is a connected linear algebraic group over $\mathbb{K}$ that acts algebraically on $A$. We identify the coordinate ring $\mathbb{K}[G\times A]$ with $\mathbb{K}[G]\otimes\mathbb{K}[A]$ where $\otimes$ denotes the tensor product over $\mathbb{K}$. The action of $G$ on $A$ determines a map
@@
For any ideal $I\subseteq\mathbb{K}[A]$ define $V(I):=\{a\in A:f(a)=0\text{ for all }f\in I\}$ and for any subset $V\subseteq A$ define $I(V):=\{f\in\mathbb{K}[A]:f(v)=0\text{ for all }v\in V\}.$ Because $\mathbb{K}$ is algebraically closed, if $I$ is any ideal then $V(I)$ is a closed affine subvariety with $I(V(I))=\sqrt{I}$ equal to the radical of $I$.
The ideal $\mathrm{cl}_{G}(I)$ is related to the closure of the union of the $B$-orbits through all points in $V(I)$.
@@
Let $\pi:G\times A\to A$ be projection onto the second factor $(g,a)\mapsto a$. Then $\pi^{*}(f)=1\otimes f$ so $\pi^{*}\circ\iota^{*}(f_{1}\otimes f_{2})=f_{1}(1)\otimes f_{2}\in\mathbb{K}[G]\otimes\mathbb{K}[A]$ for $f_{1}\in\mathbb{K}[G]$ and $f_{2}\in\mathbb{K}[A].$ Thus $\pi^{*}\circ\iota^{*}$ evaluates the first factor of an element of $\mathbb{K}[G]\otimes\mathbb{K}[A]$ at $1\in G$, so this composition restricts for any ideal $I\subseteq\mathbb{K}[A]$ to a surjective $\mathbb{K}$-linear map $\mathbb{K}[G]\otimes I\to\mathbb{K}\otimes I$.
@@ 3.4 Reductions
We prove two technical lemmas in this section in order to efficiently make these reductions later. Continue to let $G$ be a linear algebraic group that acts algebraically on an affine variety $A$, with $G$ and $A$ both defined over any algebraically closed field $\mathbb{K}$. Recall that we identify $\mathbb{K}[A]$ with the polynomial ring $\mathbb{K}[x_{i}:i\in\mathsf{Index}]$ where $\mathsf{Index}$ is some finite indexing set.
@@
Our second lemma is a straightforward property of maximal ideals in polynomials rings.
@@ 4 Matrix Schubert varieties
$\mathsf{MSC}_{w}$ =$=\left\{M\in\mathsf{Mat}_{m\times n}:\text{$\operatorname{rank}M_{[i][j]}=\operatorname{rank}w_{[i][j]}$ for $(i,j)\in[m]\times[n]$}\right\},$ (4.1)
$\mathsf{MSV}_{w}$ =$=\left\{M\in\mathsf{Mat}_{m\times n}:\text{$\operatorname{rank}M_{[i][j]}\leq\operatorname{rank}w_{[i][j]}$ for $(i,j)\in[m]\times[n]$}\right\},$
@@
To study matrix Schubert varieties, one is naturally lead to investigate the vanishing ideals $I(\mathsf{MSV}_{w})$. From this perspective, it is of interest to compute the initial ideal of $I(\mathsf{MSV}_{w})$ and then to construct a corresponding Gröbner basis. We explain in this section how transition systems provide a streamlined way of doing such calculations. This will recover several results from [23].
The transition system approach is best suited to term orders on $\mathbb{K}[\mathsf{Mat}_{m\times n}]$ with a certain antidiagonal property. For simplicity, we will always work with the following instance of such an order. First, identify the coordinate ring of $\mathsf{Mat}_{m\times n}$ with $\mathbb{K}[\mathsf{Mat}_{m\times n}]=\mathbb{K}[x_{ij}:1\leq i\leq m,1\leq j\leq n]$ by setting $x_{ij}$ to be the linear function on matrices given by $x_{ij}(M)=M_{ij}\in\mathbb{K}$. Then order the monomials in this ring using the reverse lexicographic term order explained in Remark 2.3.
@@
Prior to [23], the problem of finding Gröbner bases for various special cases of the ideals $I(\mathsf{MSV}_{w})$ had already appeared in a number of places; see [23, §2.4] for a historical overview. Since [23], there has only been partial progress on understanding the Gröbner geometry of matrix Schubert varieties for arbitrary term orders; see, for example, [14, 19, 22, 24].
@@ 4.1 A reverse lexicographic transition system
The variety $\mathsf{MSV}_{w}$ is irreducible and equal to the Zariski closure of $\mathsf{MSC}_{w}$ [30, Ch. 15]. Let $\mathsf{B}_{m}$ and $\mathsf{B}_{n}$ be the Borel subgroups of invertible lower-triangular matrices in $\mathsf{Mat}_{m\times m}$ and $\mathsf{Mat}_{n\times n}$. Then $\mathsf{B}_{m}\times\mathsf{B}_{n}$ acts algebraically on $\mathsf{Mat}_{m\times n}$ by $(g,h)\cdot M=gMh^{\top}$, and $\mathsf{MSC}_{w}$ is the orbit of $w$ under this action. It follows that in the notation of Section 3.3, we can express
@@ 4.2 Finite indexing set
Observe that $\mathsf{MSV}_{w}=\mathsf{MSV}_{gwh^{\top}}$, $I_{w}=I_{gwh^{\top}}$, and $J_{w}=J_{gwh^{\top}}$ for all $(g,h)\in\mathsf{B}_{m}\times\mathsf{B}_{n}$ and $w\in\mathsf{Mat}_{m\times n}$. Meanwhile, the $\mathsf{B}_{m}\times\mathsf{B}_{n}$-orbit of any element of $\mathsf{Mat}_{m\times n}$ contains a unique $m\times n$ partial permutation matrix [30, Ch. 15], so the size of $\mathcal{T}$ in Theorem 4.3 is at most the number of such matrices. The following notation is convenient to exploit these observations.
@@
is in bijection with the finite set of $m\times n$ partial permutation matrices, via the map that identifies $w\in S^{(m,n)}_{\infty}$ with the $m\times n$ matrix having $1$ in position $(i,w(i))$ for each $i\in[m]$ and 00 elsewhere. We freely make this identification in order to define $\mathsf{MSV}_{w}$, $I_{w}$, and $J_{w}$ for $w\in S^{(m,n)}_{\infty}$. Then
@@
for $\lambda=(3,2)$. The outer corners of $\mathsf{dom}(w)$ are therefore $(1,4)$, $(2,3)$, and $(3,1)$.
@@
Now suppose $(i,j)\notin\mathsf{dom}(w)$. Then $\operatorname{rank}w_{[k][l]}\geq\operatorname{rank}w_{[i][j]}\geq 1$ whenever $m\geq k\geq i$ and $n\geq l\geq j$. It follows that $\mathsf{MSV}_{w}$ contains the matrix $E_{ij}$ with $1$ in position $(i,j)$ and 00 in all other positions, so $x_{ij}\notin I(\mathsf{MSV}_{w})$ as $x_{ij}(E_{ij})=1\neq 0$. It also follows that $x_{ij}$ does not occur as any of the monomials $\operatorname{adiag}(\mathcal{X}_{RC})$ generating $J_{w}$ in (4.4), so $x_{ij}\notin J_{w}$ since $J_{w}$ is a monomial ideal. ∎
For positive integers $i<j$ let $t_{ij}=(i,j)\in S_{\infty}$ be the transposition swapping $i$ and $j$. We write $u\lessdot v$ for $u,v\in S_{\infty}$ if $v$ covers $u$ in the Bruhat order on $S_{\infty}$. Recall that this holds precisely when $v=ut_{ij}$ for integers $i<j$ with $u(i)<u(j)$ such that no $i<e<j$ has $u(i)<u(e)<u(j)$.
Suppose $0_{m\times n}\neq w\in S_{\infty}^{(m,n)}$ so that $I(\mathsf{MSV}_{w})$ is not a maximal ideal. Then there exists an outer corner of $\mathsf{dom}(w)$ in $[m]\times[n]$. Choose any such outer corner $(p,q)\in[m]\times[n]$ and set
@@
It is known that $\varnothing\subsetneq\mathcal{C}(w)\subseteq S_{\infty}^{(m,n)}$ by [29, Lem. 5.8]. Finally define
@@
To make these definitions canonical, one can take $(p,q)$ to be the lexicographically minimal outer corner of $\mathsf{dom}(w)$ in $[m]\times[n]$, although any outer corner works just as well.
@@
For the other axioms, fix $w\in S_{\infty}^{(m,n)}$ with $I(\mathsf{MSV}_{w})$ not maximal. Choose an outer corner $(p,q)\in[m]\times[n]$ of $\mathsf{dom}(w)$, and define $f=x_{pq}\notin I(\mathsf{MSV}_{w})$ and $\Phi\subseteq\mathcal{T}$ as in (4.8) and (4.10). The identities required for axiom (T3) when $(I,J)=(I(\mathsf{MSV}_{w}),J_{w})$ are provided in [29, Lem. 5.11].
@@
Corollary 3.6 and Theorem 4.3 imply that the generating set in (4.3) is a Gröbner basis for $I(\mathsf{MSV}_{w})$. Thus $I_{w}=I(\mathsf{MSV}_{w})$ is prime as $\mathsf{MSV}_{w}$ is irreducible [30, Ch. 15], and $J_{w}=\operatorname{init}(I_{w})$. ∎

### v3

@@ 3.2 Examples
In these examples, $A$ will be some affine space of matrices over $\mathbb{K}$. Given $w\in A$, we let
@@
Our first example, while particularly simple, helps illustrate the general setup. Suppose $A$ is the variety of skew-symmetric $3\times 3$ matrices
@@
$J_{k}=\operatorname{init}(I_{k})\quad\text{and}\quad\mathcal{T}=\{(I_{k},J_{k}):k=1,2,3,4,5,6,7\}.$
We now investigate whether $\mathcal{T}$ is a transition system. As one can informally express
$X_{1}$ $=\left\{\begin{bmatrix}0&0\\ 0&0\end{bmatrix}\right\},\quad X_{2}=\left\{\begin{bmatrix}0&0\\ 0&\ast\end{bmatrix}\right\},\quad X_{3}=\left\{\begin{bmatrix}0&0\\ \ast&\ast\end{bmatrix}\right\},\quad X_{4}=\left\{\begin{bmatrix}0&\ast\\ 0&\ast\end{bmatrix}\right\}$,
$X_{5}$ $=\left\{\begin{bmatrix}0&\ast\\ \ast&\ast\end{bmatrix}\right\},\quad X_{6}=\left\{\begin{bmatrix}\ast&\ast\\ \ast&\ast\end{bmatrix}\right\},\quad X_{7}=\left\{\det=0\right\},$
it follows that
@@
We only need to verify conditions (T3) and (T4) in Definition 3.1 when $(I,J)=(I_{7},J_{7})$. First suppose $J_{7}=\langle x_{12}x_{21}\rangle$. Then
$I_{7}+\langle x_{11}\rangle=J_{7}+\langle x_{11}\rangle=\langle x_{11},x_{12}x_{21}\rangle=I_{3}\cap I_{4}=J_{3}\cap J_{4},$
so $f=x_{11}$ and $\Phi=\{(I_{3},J_{3}),(I_{4},J_{4})\}$ satisfy condition (T3) in Definition 3.1. For these choices, condition (T4) holds vacuously, and we conclude that $\mathcal{T}$ is a transition system.
@@
The set $\{(I_{k},J_{k}):k=1,2,3,4,5\}$ fails to be a transition system, though only by a small margin.
We only need to verify conditions (T3) and (T4) in Definition 3.1 when $(I,J)=(I_{1},J_{1})$ since $I_{k}=J_{k}$ when $k>1$. The problem is that $J_{1}+\langle\operatorname{init}(f)\rangle$ is never equal to $\bigcap_{k\in S}J_{k}$ for any choice of set $\varnothing\subsetneq S\subseteq\{1,2,3,4,5\}$ when $f\in\mathbb{K}[A]$ is non-constant with $f\notin I_{1}$. To get around this, define
@@ 3.3 Orbit closures
We are most interested in transition systems $\mathcal{T}$ made up of pairs $(I,J)$ where $I$ ranges over the orbit closures of certain maximal ideals. This section studies the relevant orbit closure operation. It is convenient to present these results in an entirely self-contained way, but generalizations of this material in terms of equivariant sheaves may be well-known; see, e.g., [5, 32]. If the reader is willing to believe that obvious properties of orbit closures of subvarieties extend naturally from radical ideals to all ideals, the details of this section can be skipped.
Continue to let $A$ be an affine space over an algebraically closed field $\mathbb{K}$. Suppose $G$ is a connected linear algebraic group over $\mathbb{K}$ that acts algebraically on $A$. We identify the coordinate ring $\mathbb{K}[G\times A]$ with $\mathbb{K}[G]\otimes\mathbb{K}[A]$ where $\otimes$ denotes the tensor product over $\mathbb{K}$. The action of $G$ on $A$ determines a map
@@
For any ideal $I\subseteq\mathbb{K}[A]$ define $V(I):=\{a\in A:f(a)=0\text{ for all }f\in I\}$ and for any subset $V\subseteq A$ define $I(V):=\{f\in\mathbb{K}[A]:f(v)=0\text{ for all }v\in V\}.$ Because $\mathbb{K}$ is algebraically closed, Hilbert’s Nullstellensatz asserts that if $I$ is any ideal then $V(I)$ is a closed affine subvariety and $I(V(I))=\sqrt{I}$ is equal to the radical of $I$.
The ideal $\mathrm{cl}_{G}(I)$ is related to the closure of the union of the $G$-orbits through all points in $V(I)$.
@@
Let $\pi:G\times A\to A$ be the projection map onto the second factor $(g,a)\mapsto a$. Then $\pi^{*}(f)=1\otimes f$ so $\pi^{*}\circ\iota^{*}(f_{1}\otimes f_{2})=f_{1}(1)\otimes f_{2}\in\mathbb{K}[G]\otimes\mathbb{K}[A]$ for $f_{1}\in\mathbb{K}[G]$ and $f_{2}\in\mathbb{K}[A].$ Thus $\pi^{*}\circ\iota^{*}$ evaluates the first factor of an element of $\mathbb{K}[G]\otimes\mathbb{K}[A]$ at $1\in G$, so this composition restricts for any ideal $I\subseteq\mathbb{K}[A]$ to a surjective $\mathbb{K}$-linear map $\mathbb{K}[G]\otimes I\to\mathbb{K}\otimes I$.
@@ 3.4 Reductions
We prove two technical lemmas in this section in order to efficiently make these reductions later. Continue to let $G$ be a linear algebraic group that acts algebraically on an affine space $A$, with $G$ and $A$ both defined over any algebraically closed field $\mathbb{K}$. Recall that we identify $\mathbb{K}[A]$ with the polynomial ring $\mathbb{K}[x_{i}:i\in\mathsf{Index}]$ where $\mathsf{Index}$ is some finite indexing set.
@@
Our second lemma is a straightforward property of maximal ideals in polynomial rings.
@@ 4 Matrix Schubert varieties
$\mathsf{MSC}_{w}$ $=\left\{M\in\mathsf{Mat}_{m\times n}:\text{$\operatorname{rank}M_{[i][j]}=\operatorname{rank}w_{[i][j]}$ for $(i,j)\in[m]\times[n]$}\right\},$ (4.1)
$\mathsf{MSV}_{w}$ $=\left\{M\in\mathsf{Mat}_{m\times n}:\text{$\operatorname{rank}M_{[i][j]}\leq\operatorname{rank}w_{[i][j]}$ for $(i,j)\in[m]\times[n]$}\right\},$
@@
To study matrix Schubert varieties, one is naturally led to investigate the vanishing ideals $I(\mathsf{MSV}_{w})$. From this perspective, it is of interest to compute the initial ideal of $I(\mathsf{MSV}_{w})$ and then to construct a corresponding Gröbner basis. We explain below how to do such calculations using transition systems. This will recover several results from [25].
The transition system approach is best suited to term orders on $\mathbb{K}[\mathsf{Mat}_{m\times n}]$ with a certain antidiagonal property defined in [25, §1.5]. For simplicity, we will always work with the following instance of such an order. First, identify the coordinate ring of $\mathsf{Mat}_{m\times n}$ with
$\mathbb{K}[\mathsf{Mat}_{m\times n}]=\mathbb{K}[x_{ij}:1\leq i\leq m,1\leq j\leq n]$
by setting $x_{ij}$ to be the linear function on matrices given by $x_{ij}(M)=M_{ij}\in\mathbb{K}$. Then order the monomials in this ring using the reverse lexicographic term order explained in Remark 2.3.
@@
Prior to [25], the problem of finding Gröbner bases for various special cases of the ideals $I(\mathsf{MSV}_{w})$ had already appeared in a number of places; see [25, §2.4] for a historical overview. Since [25], there has only been partial progress on understanding the Gröbner geometry of matrix Schubert varieties for arbitrary term orders; see, for example, [14, 19, 24, 26].
@@ 4.1 A reverse lexicographic transition system
The variety $\mathsf{MSV}_{w}$ is irreducible and equal to the Zariski closure of $\mathsf{MSC}_{w}$ [32, Ch. 15]. Let $\mathsf{B}_{m}$ and $\mathsf{B}_{n}$ be the Borel subgroups of invertible lower-triangular matrices in $\mathsf{Mat}_{m\times m}$ and $\mathsf{Mat}_{n\times n}$. Then $\mathsf{B}_{m}\times\mathsf{B}_{n}$ acts algebraically on $\mathsf{Mat}_{m\times n}$ by $(g,h)\cdot M=gMh^{\top}$, and $\mathsf{MSC}_{w}$ is the orbit of $w$ under this action. It follows that in the notation of Section 3.3, we can express
@@ 4.2 Finite indexing set
Observe that $\mathsf{MSV}_{w}=\mathsf{MSV}_{gwh^{\top}}$, $I_{w}=I_{gwh^{\top}}$, and $J_{w}=J_{gwh^{\top}}$ for all $(g,h)\in\mathsf{B}_{m}\times\mathsf{B}_{n}$ and $w\in\mathsf{Mat}_{m\times n}$. Meanwhile, the $\mathsf{B}_{m}\times\mathsf{B}_{n}$-orbit of any element of $\mathsf{Mat}_{m\times n}$ contains a unique $m\times n$ partial permutation matrix [32, Ch. 15], so the size of $\mathcal{T}$ in Theorem 4.3 is at most the number of such matrices. The following notation is convenient to exploit these observations.
@@
is in bijection with the finite set of $m\times n$ partial permutation matrices, via the map that identifies $w\in S^{(m,n)}_{\infty}$ with the $m\times n$ matrix having $1$ in position $(i,w(i))$ for each $i\in[m]$ and 0 elsewhere. We freely make this identification in order to define $\mathsf{MSV}_{w}$, $I_{w}$, and $J_{w}$ for $w\in S^{(m,n)}_{\infty}$. Then
@@
where we compute $\operatorname{rank}w_{[i][j]}$ by identifying $w$ with the (infinite) permutation matrix having $1$ in each position $(i,w(i))$. The set $\mathsf{dom}(w)$ always coincides with the Young diagram
[…]
@@
for $\lambda=(3,2)$. The outer corners of $\mathsf{dom}(w)$ are $(1,4)$, $(2,3)$, and $(3,1)$.
@@
Now suppose $(i,j)\notin\mathsf{dom}(w)$. Then $\operatorname{rank}w_{[k][l]}\geq\operatorname{rank}w_{[i][j]}\geq 1$ whenever $m\geq k\geq i$ and $n\geq l\geq j$. It follows that $\mathsf{MSV}_{w}$ contains the matrix $E_{ij}$ with $1$ in position $(i,j)$ and 0 in all other positions, so $x_{ij}\notin I(\mathsf{MSV}_{w})$ as $x_{ij}(E_{ij})=1\neq 0$. It also follows that $x_{ij}$ does not occur as any of the monomials $\operatorname{adiag}(\mathcal{X}_{RC})$ generating $J_{w}$ in (4.4), so $x_{ij}\notin J_{w}$ since $J_{w}$ is a monomial ideal. ∎
For positive integers $i<j$ let $t_{ij}=(i,j)\in S_{\infty}$ be the transposition swapping $i$ and $j$. We write $u\lessdot v$ for $u,v\in S_{\infty}$ if $v$ covers $u$ in the Bruhat order on $S_{\infty}$. Recall from [2, Lem. 2.1.4] that this holds precisely when $v=ut_{ij}$ for integers $i<j$ with $u(i)<u(j)$ such that no $i<e<j$ has $u(i)<u(e)<u(j)$.
Suppose $0_{m\times n}\neq w\in S_{\infty}^{(m,n)}$ so that $I(\mathsf{MSV}_{w})$ is not a maximal ideal. Then there exists an outer corner of $\mathsf{dom}(w)$ in $m]\times[n]$. Choose any such outer corner $(p,q)\in[m]\times[n]$ and set
@@
It is known that $\varnothing\subsetneq\mathcal{C}(w)\subseteq S_{\infty}^{(m,n)}$ by [31, Lem. 5.8]. Finally define
@@
To make these definitions canonical, one can take $(p,q)$ to be the lexicographically minimal outer corner of $\mathsf{dom}(w)$ in $m]\times[n]$, although any outer corner works just as well.
@@
For the other axioms, fix $w\in S_{\infty}^{(m,n)}$ with $I(\mathsf{MSV}_{w})$ not maximal. Choose an outer corner $(p,q)\in[m]\times[n]$ of $\mathsf{dom}(w)$, and define $f=x_{pq}\notin I(\mathsf{MSV}_{w})$ and $\Phi\subseteq\mathcal{T}$ as in (4.9) and (4.11). The identities required for axiom (T3) when $(I,J)=(I(\mathsf{MSV}_{w}),J_{w})$ are provided in [31, Lem. 5.11].
@@
Corollary 3.6 and Theorem 4.3 imply that the generating set in (4.3) is a Gröbner basis for $I(\mathsf{MSV}_{w})$. Thus $I_{w}=I(\mathsf{MSV}_{w})$ is prime as $\mathsf{MSV}_{w}$ is irreducible [32, Ch. 15], and $J_{w}=\operatorname{init}(I_{w})$. ∎